1.A candy manufacturer has two grades of candy, selling at 75 cents and 90 cents a pound respectively. How many pounds of each must he use to make mixture of 150 pounds, to sell at 75 cents a pound? please, show your step by step solution..
2. A tobacconist wishes to blend two grades of tobacco, costing 8 cents an ounce and 10 cents an ounce respectively, and to sell the mixture at 12 cents an ounce with a profit of 25 per cent of the cost.How many ounces of each kind must he used to the pound ? ( 1 lb. = 16 oz.)
3. A fruit dealer sold 9 dozen lemons and 16 dozen oranges for $17.40.The following day he sold, at the same prices, 11 dozens lemons and 20 dozens oranges for $21.60. Find the prices of his lemons and oranges.

Answer by mananth(16949)

----------------- price cents ----------------quantity
Candy type I 75 ----------------x pounds
Candy type II 90 ----------------600-x pounds
Mixture 75.00 ----------------600

75x+90(600-x)=75*600

75x+54000-90x=45000
75x-90x=45000-54000
-15x=-9000
/-15
x=600 pounds 75.00 cents Candy type I
0 pounds 90.00 cents Candy type II



tobacco type I - 8 cents/ounce x ounces
tobacco type II- 10 cents/ounce y ounces
he sells at 12 cents/pound
profit = 25%
so cost = (100/125)*12=9.60 cents
sum of individual costs = costof mixture
8x+10y= 9.6(x+y)
8x+10y=9.6x+9.6y
0.4y=1.6x
x/y = 0.4/1.6
x/y = 1/4
he has to mix 3.2 ounces of 8 cents cent tobacco with 12.8 ounce of 10 cent tobacco to make 1 pound of mixed coffee
m.ananth@hotmail.ca